OpenStudy (anonymous):

what is the slope of the line that passes through the origin and the point (5,2) ? a) m=-2.5 b)m=-2/5 c)m=4 d)m=0.4

OpenStudy (anonymous):

here when the say origin they are giving you a second point (0,0). now you have two points. finding the slope should be easy. if not then type "slope khan academy" into google.

OpenStudy (anonymous):

@tomo so what you think the answer is ?

OpenStudy (anonymous):

hmm, that would be too easy.

OpenStudy (anonymous):

just type slope formula into google.

OpenStudy (anonymous):

@tomo thats my problem, my computer isn't letting me !

OpenStudy (anonymous):

ok, here is the formula. \[m = \frac{y2-y1}{x2-x1}\] m is the slope (x1,y1) = (0,0) (x2,y2) = (5,2) what do you get as the value for m

OpenStudy (anonymous):

@tomo i got 1

OpenStudy (anonymous):

@tomo actually 2/5